# A Decision Theoretic Framework for Measuring AI Reliance

### v4

@@ 1. Introduction
AI-advised decision making, in which a human decision-maker has access to the recommendation of an artificial intelligence (AI system) and can choose whether or not to follow it, is often preferred as a means of retaining human control (Bansal et al., 2021) in deploying predictive models. The motivation behind this approach is complementary performance; i.e., the human-AI team can outperform the AI or the human alone. However, many studies have shown that human-AI teams under-perform the AI alone in tasks where the AI’s accuracy is higher than humans (Bansal et al., 2021; Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). One solution to this problem is to identify ways to ensure that the human, as the final decision-maker, has appropriate reliance on AI. Appropriate reliance is typically defined as submitting the AI recommendation when it is correct and not submitting it when it is not correct.
@@
Implicit in discussions of complementarity are assumptions of a human with some internal model of the data-generating process and an AI with its own model. Studying reliance implies that the human consults the AI recommendation, infers the probability that its decision is correct, then decides whether it is worth following its recommendation. Problems arise because defining appropriate reliance as submitting the AI’s recommendation when it is correct and rejecting it when it is not confounds two challenges a human may face in an AI-advised decision-making: that of forming correct beliefs about the probability that the AI is correct, and that of making the optimal decision about whether to follow the AI conditional on one’s beliefs. Without a definition that allows separation of different sources of performance loss, the analysis might misinterpret the reasons behind seeminly poor experiment results, leading researchers to prioritize less directly relevant follow-up actions for improving the team. For example, if the human has inaccurate beliefs about the probability that the AI is correct, this might stem from a lack of information about the prior probability that the AI is correct (potentially addressable by providing the AI’s accuracy on held-out data (Yin et al., 2019)), or from their failure to arrive at an accurate estimate of the AI’s probability of being correct (potentially fixable via cognitive forcing functions (Buçinca et al., 2021; Fogliato et al., 2021) or better explanations (Bansal et al., 2021)). If the human correctly perceives the accuracy of the AI model, but uses the wrong decision rule to decide when to follow its recommendation, then the human may not understand the utility of different possible outcomes (e.g., a differential cost of using the AI’s recommendation versus generating their own), or the researcher studying real-world human-AI teams may have assumed a utility function different from that used by the participant.
@@
We propose a formal definition of AI reliance. Following previous work on generating benchmarks for studies of information displays (Wu et al., 2023), our approach is grounded in statistical decision theory. Our definition separates the concepts of a reliance level (the probability that the human decision-maker goes with the AI recommendation) from the belief updating that a rational decision-maker is expected to do upon viewing an instance and associated AI recommendation. The framework we provide defines a benchmark for complementary performance representing the maximum attainable performance with the cooperation of AI and human and a baseline for complementary performance representing the maximum performance without any cooperation. We apply the framework to three well-regarded AI-advised decision making experiments from literature (Bansal et al., 2021; Lai and Tan, 2019; Fogliato et al., 2021). In all three cases, we show 1) that examining the results against the baseline and benchmark for complementary performance better reveals the limits of human behavioral performance and 2) specific sources of behavioral loss that help explain the experiment results but were not accounted for by the original interpretations of the results.
@@ 2. Formulating Assumptions for Studying Reliance
    The human makes their own prediction about each instance prior to seeing the AI recommendation for that instance.
@@
There are two benefits to making these assumptions for AI-advised decision-making experiments. First, the assumptions ensure that participants neither anchor solely on the AI recommendations (completely neglecting to consider their own predictions) nor that they neglect to consult the AI recommendation at all (Buçinca et al., 2021; Fogliato et al., 2021). It is difficult to conceive of reliance in such cases. Second, and most importantly, by assuming we have access to the human’s own prediction prior to their interaction with the AI recommendation, we can compare the results of experiments we run to a benchmark of complementary performance, which is attained by optimally combining the information contained in the human’s predictions with that contained in the AI’s recommendations, and a baseline of using either the AI or human only. We use human recommendation to refer to the human prediction prior to interaction with the AI recommendation.
@@ 3. Definition of Reliance
We define appropriate reliance, over-reliance, and under-reliance on AI recommendations in AI-advised decision making. Our framework conceives of three roles in the decision problem: a human recommender, an AI recommender, and a decision-maker. The two recommenders provide informational input to the decision-maker in the form of recommendations. The decision-maker chooses which recommender to follow on a decision task.
To formalize a decision task requires five key elements (Table 1): payoff-related states on which the decision is evaluated, a data genertaing model that generates the states and signals that inform about the state, the action, the information (i.e. signal) given to the decision-maker, and a scoring rule assessing the choice of action under the payoff-related state.
@@
Explanation $e(y^{AI})$
Action (choice) $a\in Y$ $\hat{a}\in\{0=\text{human},1=\text{AI}\}$
Signal $v=\{x,y^{H},y^{AI},e(y^{AI})\}$
@@ 3.1. Rational Decision-Maker
We define the rational decision-maker in a binary-adoption decision task (Table 1) derived from the original one. This derived decision task limits the rational decision-maker to making a final decision by selecting between the human recommendation and the AI recommendation. We define the rational benchmark representing the expected performance of a rational Bayesian decision-maker who perfectly perceives the provided information in the signal and chooses the optimal action under the scoring rule for each decision task. The rational benchmark is the maximum payoff that can be expected from a behavioral decision-maker, i.e., the benchmark for complementary performance. Following the framework proposed by Wu et al. (Wu et al., 2023), we also define a baseline for expected performance using this rational Bayesian decision-maker. The rational baseline is the maximum payoff that can be expected from the behavioral decision-maker when they must choose between always going with either the AI or the human recommender, i.e., they do not consult the individual signals in making their decisions. The rational baseline represents the minimum threshold for achieving complementary performance, i.e., the baseline for complementary performance. Using the rational benchmark and the rational baseline, we define the value of rational complementation, representing the expected improvement in payoff to a rational decision-maker that the joint human+AI setting provides over the better of either the AI or the human alone.
@@
  • Rational Benchmark, The rational benchmark is the expected performance of the rational decision-maker with the signal on a randomly chosen decision task from the experiment. Let $a^{r}(v)$ be the action taken by the rational decision-maker given signal $v$. She chooses $a^{r}$ to maximize her expected utility with $\pi(\hat{\theta}|v)$, the distribution of the payoff-related state conditioned on the signal $v$:
    $\textsc{R}=\max_{a^{r}(\cdot)}\text{\bf E}_{\pi(v,\hat{\theta})}\![{\hat{S}(a^{r}(v),\hat{\theta})}].$
@@
The value of rational complementation provides a scale for comparing expected performance in terms of the “lift” we see from having access to the information in the signals. In the context of AI-advised decision making, it also represents the maximum improvement of performance we can expect from a complementation of the human and the AI conditioned on the information structure of the signals. If we treat $\Delta$ as a comparative unit by normalizing all scores within the range where the baseline $\textsc{R}_{\varnothing}$ is 00 and the benchmark R is $1$, we get a sense of the proportion of possible score increase that different settings provide. For example, we could compare expected human performances $\textsc{B}_{\alpha}$ and $\textsc{B}_{\beta}$ under two conditions $\alpha$ and $\beta$ (e.g., $\alpha$ explanation and $\beta$ explanation) by calculating $(\textsc{B}_{\alpha}-\textsc{B}_{\beta})/\Delta$.
@@ 3.2. Behavioral Decision-Maker
    $\textsc{R}^{\text{m}}(\gamma)=\max_{a^{r}(\cdot)}\quad$ $\text{\bf E}_{\pi(\hat{v},\hat{\theta})}\![{\hat{S}(a^{r}(\cdot),\hat{\theta})}]$
@@ 4. Applying the Framework to AI Reliance Studies
Experiment design and data collection. The first step in applying the framework is to formulate the experiment design as a decision problem by defining the ground truth state, data-generating model, action space, signal, and scoring rule. Imagine we run an experiment studying AI-advised recidivism decisions with $200$ humans, where each completes 20 trials. In each trial they view a profile of the defendent, and must predict whether the defendent will be re-arrested. The participants are assisted with an AI model that is deterministic and calibrated on the ground truth. We equally divide the $200$ participants into two groups, randomly assigning $100$ to one explanation condition and the other $100$ to a different explanation condition. All participants first do the $20$ instances by themselves before they see any AI recommendations, then make final decisions on the same $20$ instances with the AI assistance. For every correct decision on the second batch of trials, the participant receives \$$\$0.5$ as incentivization. The decision tasks are formalized in Table 2 in Appendix B. When the experiment is complete, we have collected $4000$ decision observations in total. Each observation includes information about the profile of the defendent, the outcome of whether the defendent is re-arrested, the human recommendation on the first batch of trials, the AI recommendation, the explanation of the AI recommendation, and the final decision on the second batch of trials.
@@
(Approximating) Rational benchmark R. To calculate the rational benchmark we identify the best response to each signal. When the signal space has finite size, we can calculate the rational benchmark by simulating the best response to each signal on the empirical distribution of the experiment observations. However, for a large number of decision tasks in the literature (including, e.g., the demonstrations in Section 5), the signal space has near infinite size (e.g., it involves text documents) such that each experimental observation might involve a different unique signal. Thus, the identified best response action may overfit to the data relative to the true expected score of the rational decision-maker on a randomly chosen decision task from the experiment. We approximate the rational benchmark by designing an upperbound and a lowerbound.
  • Upperbound: Overfitting to the empirical distribution. We calculate the rational benchmark on the empirical joint distribution $\tilde{\pi}(\hat{\theta},v)$ over the payoff-relevant state $\hat{\theta}$ and the signal $v$, treating the empirical distribution as the true data generating model. Algorithm 2 in Appendix A calculates this empirical distribution.
To see why this is an upperbound and why we call it overfitting, consider the case where the signal space is continuous. Each entry in the experiment data has a distinct signal. Without repetition, it is impossible to approximate the true distribution of the payoff-relevant state $\hat{\theta}$ conditioning on each signal $v$. Treating the empirical distribution as the true data generating model, there is no randomness in the payoff-relevant state given the rational decision-maker’s knowledge.
  • Lowerbound: Learning the best response on the optimally discretized empirical distribution to avoid overfitting. Assuming continuity on the joint distribution $\tilde{\pi}(\hat{\theta},v)$ over the payoff-relevant state $\hat{\theta}$ and the signal $v$, we approximate the rational benchmark by coarsening the signal space into finite discrete signals $\tilde{v}_{1},\tilde{v}_{2},\dots,\tilde{v}_{k}$, and calculating the best response on the empirical distribution over the discretized space $\{\tilde{v}_{i}\}_{i}$. An example using the $k$-means algorithm to discretize the signals is shown in Algorithm 3 in Appendix A.
@@
The calculation of the rational benchmark hence takes an empirical distribution as input. For a finite signal space, the rational benchmark is calculated on the empirical distribution. For an infinite signal space, the upperbound is calculated on the empirical distribution, while the lowerbound is calculated on the discretized empirical distribution. Regardless of which bound we are calculating, given an empirical distribution (e.g, the $4000$ observations), we simulate the rational decision-maker’s decision. For each observation, the rational decision-maker receives a signal (raw signal or discretized signal) and calculates the posterior distribution of states given the signal by Bayes rule, denoted as $\pi(\hat{\theta}|v)=\frac{\pi(\hat{\theta},v)}{\pi(v)}$. We pick the action with higher expected payoff under the posterior distribution on the current observation. We repeat this process for all observations and then take the expectation on all the rational benchmarks we get. We can take the conditional expectation across different conditions, e.g., different explanations. This calculation is illustrated in Algorithm 4 in Appendix A.
@@
(Approximating) Mis-reliant rational benchmark $\textsc{R}^{\text{m}}$. The mis-reliant rational benchmark is the expected score of a rational decision-maker with the same behavioral reliance level as the human participant. To calculate this, we simulate the rational decision-maker completing the same set of trials as the behavioral decision-makers do but additionally constrain the reliance level to be the same as the reliance level produced by the behavioral decision-makers. In our example experiment, each behavioral decision-maker completes $20$ trials with reliance levels, $\gamma^{b}=\Pr[a^{b}=y^{AI}|y^{AI}\neq y^{H}]$. As the rational decision-maker traverses the $4000$ observations, like behavioral participants she should engage in $20$ consecutive trials for each set. Suppose that the signals that the rational decision-maker receives in the $20$ consecutive trials are $v_{1},\ldots,v_{20}$. For each signal $v_{i}$, the rational decision-maker knows the posterior payoffs, i.e., $\text{\bf E}_{\pi(\hat{\theta}|v_{i})}\![{S(y^{AI},y)}]$ and $\text{\bf E}_{\pi(\hat{\theta}|v_{i})}\![{S(y^{H},y)}]$. Then, the rational decision-maker ranks the signals in decreasing order of $\text{\bf E}_{\pi(\hat{\theta}|v_{i})}\![{S(y^{AI},y)}]-\text{\bf E}_{\pi(\hat{\theta}|v_{i})}\![{S(y^{H},y)}]$ and accepts the AI recommendation from the first signal in the sorted list, up to a $\gamma^{b}$ fraction of $20$ signals. We take the expectation over all observations (or conditionally on the manipulated variable of interest depending on the study design). This calculation is illustrated in Algorithm 6 in Appendix A. Note that estimation of the mis-reliant rational benchmark faces the same risk of overfitting as the rational benchmark. When the signal space is infinite, we approximate the mis-reliant rational benchmark the same way that we do the rational benchmark by calculating the upper- and lower-bound.

### v5

@@ 1. Introduction
AI-advised decision making, in which a human decision-maker has access to the recommendation of an artificial intelligence (AI system) and can choose whether or not to follow it, is often preferred as a means of retaining human control (Bansal et al., 2021) in deploying predictive models. The motivation behind this approach is complementary performance; i.e., the human-AI team can outperform the AI or the human alone. However, many studies have shown that human-AI teams under-perform the AI alone in tasks (Bansal et al., 2021; Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). One solution to this problem is to identify ways to ensure that the human, as the final decision-maker, has appropriate reliance on AI, which is typically defined as submitting the AI recommendation when it is correct and not submitting it when it is not correct.
@@
Implicit in discussions of complementarity are assumptions of a human with some internal model of the data-generating process and an AI with its own model. Studying reliance implies that the human consults the AI recommendation, infers the probability that its decision is correct, then decides whether it is worth following its recommendation. Problems arise because defining appropriate reliance as submitting the AI’s recommendation when it is correct and rejecting it when it is not confounds two challenges a human may face in an AI-advised decision-making: that of forming correct beliefs about the probability that the AI is correct (i.e., prior beliefs), and that of making the optimal decision about whether to follow the AI conditional on the observed information (i.e., posterior beliefs). Without a definition that allows separation of different sources of performance loss, the analysis might misinterpret the reasons behind seemingly poor experiment results, leading researchers to prioritize less directly relevant follow-up actions for improving the team. For example, if the human has inaccurate beliefs about the probability that the AI is correct, this might stem from a lack of information about the prior probability that the AI is correct (potentially addressable by providing the AI’s accuracy on held-out data (Yin et al., 2019)). If the human correctly forms the prior belief on AI’s accuracy, but fails to correctly choose the AI’s recommendation instead of following their own decisions, then this might stem from a lack of motivation to process the information (potentially fixable via cognitive forcing functions (Buçinca et al., 2021; Fogliato et al., 2021), better design of incentivations (Li et al., 2022), or better information display such as explantions (Bansal et al., 2021)).
@@
We propose a formal definition of AI reliance. Following previous work on generating benchmarks for studies of information displays (Wu et al., 2023), our approach is grounded in statistical decision theory. Our definition separates the concepts of a reliance level (the probability that the human decision-maker goes with the AI recommendation) from the belief updating that a rational decision-maker is expected to do upon viewing an instance and associated AI recommendation. The framework evaluates these two factors of error within the payoff space (i.e.,how much gain can be achieved if the error is corrected) by introducing a benchmark for the best-attainable performance of human-AI collaboration and a baseline for the performance with no collaboration. We apply the framework to three well-regarded AI-advised decision making experiments from literature (Bansal et al., 2021; Lai and Tan, 2019; Fogliato et al., 2021). In all three cases, we show 1) that examining the results against the baseline and benchmark for complementary performance better reveals the limits of human behavioral performance and 2) specific sources of behavioral loss that help explain the experiment results but were not accounted for by the original interpretations of the results.
@@ 2. Formulating Assumptions for Studying Reliance
    The human has an inner model of the data-generating process that enables them to make their own prediction about each instance.
@@
There are two benefits to making these assumptions for AI-advised decision-making experiments. First, the assumptions ensure that participants neither anchor solely on the AI recommendations (completely neglecting to consider their own predictions) nor that they neglect to consult the AI recommendation at all (Buçinca et al., 2021; Fogliato et al., 2021). It is difficult to conceive of reliance without such assumptions. Second, and most importantly, by assuming we have access to the human’s own prediction prior to their interaction with the AI recommendation, we can compare the results of experiments we run to a benchmark of complementary performance, which is attained by optimally combining the information contained in the human’s predictions with that contained in the AI’s recommendations, and a baseline of using either the AI or human only. We use human recommendation to refer to the human prediction prior to interaction with the AI recommendation.
@@ 3. Definition of Reliance
We define appropriate reliance, over-reliance, and under-reliance on AI recommendations in AI-advised decision making. Our framework conceives of three roles in the decision problem: a human recommender, an AI recommender, and a decision-maker. The two recommenders provide informational input to the decision-maker in the form of recommendations. The decision-maker chooses which recommender to follow on a decision task after seeing the recommendations.
To formalize a decision task requires five key elements (Table˜1): payoff-related states on which the decision is evaluated, a data genertaing model that generates the states and signals that inform about the state, the action, the information (i.e. signal) given to the decision-maker, and a scoring rule assessing the choice of action under the payoff-related state.
@@
[…]
Signal $v=\{x,y^{AI}\}$ $\hat{v}=\{x,y^{H},y^{AI}\}$
@@ 3.1. Rational Decision-Maker
We define the rational decision-maker in a binary-adoption decision task (Table 1) derived from the original one. This derived decision task limits the rational decision-maker to making a final decision by selecting between the human recommendation and the AI recommendation. We define the rational benchmark representing the expected performance of a rational Bayesian decision-maker who perfectly perceives the provided information and chooses the optimal action under the scoring rule for each decision task  [^1] The expected performance of a rational decision-maker who has access to $y^{H}$, $y^{AI}$, and $x$ is an upperbound on the behavioral decision-maker. Though some explanations of $y^{AI}$ may also be available to behavioral decision-maker, they are a garbling of $x$ and $y^{AI}$ and won’t change the rational decision-maker’s action.. We use hat symbols to denote the components of the derived decision task, e.g., $\hat{\theta}$ is the payoff-related state in the derived decision task. The rational benchmark is the maximum payoff that can be expected from a behavioral decision-maker, i.e., the benchmark for complementary performance. Following the framework proposed by Wu et al. (Wu et al., 2023), we also define a baseline for expected performance using this rational Bayesian decision-maker. The rational baseline is the maximum payoff that can be expected from the behavioral decision-maker when they must choose between always going with either the AI or the human recommender, i.e., they do not consult the individual signals in making their decisions. The rational baseline represents the minimum threshold for achieving complementary performance, i.e., the baseline for complementary performance. Using the rational benchmark and the rational baseline, we define the value of rational complementation, representing the expected improvement in payoff to a rational decision-maker that the joint human+AI setting provides over the better of either the AI or the human alone.
@@
[…]
    $\textsc{R}=\max_{a^{r}(\cdot)}\text{\bf E}_{\pi(\hat{v},\hat{\theta})}\![{\hat{S}(a^{r}(\hat{v}),\hat{\theta})}$].
@@
The value of rational complementation provides a scale for comparing expected performance in terms of the “lift” we see from having access to the information in the signals. In the context of AI-advised decision making, it also represents the maximum improvement of performance we can expect from a complementation of the human and the AI conditioned on the information structure of the signals. If we treat $\Delta$ as a comparative unit by normalizing all scores within the range where the baseline $\textsc{R}_{\varnothing}$ is 0 and the benchmark R is $1$, we get a sense of the proportion of possible score increase that different settings provide. For example, we could compare expected human performances $\textsc{B}_{\alpha}$ and $\textsc{B}_{\beta}$ under two conditions $\alpha$ and $\beta$ (e.g., $\alpha$ explanation and $\beta$ explanation) by calculating $(\textsc{B}_{\alpha}-\textsc{B}_{\beta})/\Delta$.
@@ 3.2. Behavioral Decision-Maker
    $\textsc{R}^{\text{m}}(\gamma)=\max_{a^{r}(\cdot)}\quad$ $\text{\bf E}_{\pi(\hat{v},\hat{\theta})}\![{\hat{S}(a^{r}(\hat{v}),\hat{\theta})}$]
@@ 4. Applying the Framework to AI Reliance Studies
Experiment design and data collection. The first step in applying the framework is to formulate the experiment design as a decision problem by defining the ground truth state, data-generating model, action space, signal, and scoring rule. Imagine we run an experiment studying AI-advised recidivism decisions with $200$ humans, where each completes 20 trials. In each trial they view a profile of the defendent, and must predict whether the defendent will be re-arrested. The participants are assisted with an AI model that is deterministic and calibrated on the ground truth. We equally divide the $200$ participants into two groups, randomly assigning $100$ to one explanation condition and the other $100$ to a different explanation condition. All participants first do the 2020 instances by themselves before they see any AI recommendations, then make final decisions on the same $20$ instances with the AI assistance. For every correct decision on the second batch of trials, the participant receives $\mathdollar 0.5$ as incentivization. The decision tasks are formalized in Table 2 in Appendix B. When the experiment is complete, we have collected $4000$ decision observations in total. Each observation includes information about the profile of the defendent, the outcome of whether the defendent is re-arrested, the human recommendation on the first batch of trials, the AI recommendation, the explanation of the AI recommendation, and the final decision on the second batch of trials.
@@
(Approximating) Rational benchmark R. To calculate the rational benchmark we identify the best response to each signal. When the signal space has finite size, we can calculate the rational benchmark by simulating the best response to each signal on the empirical distribution of the experiment observations. However, for a large number of decision tasks in the literature (including, e.g., the demonstrations in Section˜5), the signal space has near infinite size (e.g., it involves text documents) such that each experimental observation might involve a different unique signal. Thus, the identified best response action may overfit to the data relative to the true expected score of the rational decision-maker on a randomly chosen decision task from the experiment. We approximate the rational benchmark by designing an upperbound and a lowerbound.
[…]
To see why this is an upperbound and why we call it overfitting, consider the case where the signal space is continuous. Each entry in the experiment data has a distinct signal. Without repetition, it is impossible to approximate the true distribution of the payoff-relevant state $\hat{\theta}$ conditioning on each signal $\hat{v}$. Treating the empirical distribution as the true data generating model, there is no randomness in the payoff-relevant state given the rational decision-maker’s knowledge.
[…]
@@
The calculation of the rational benchmark hence takes an empirical distribution as input. For a finite signal space, the rational benchmark is calculated on the empirical distribution. For an infinite signal space, the upperbound is calculated on the empirical distribution, while the lowerbound is calculated on the discretized empirical distribution. Regardless of which bound we are calculating, given an empirical distribution (e.g, the $4000$ observations), we simulate the rational decision-maker’s decision. For each observation, the rational decision-maker receives a signal (raw signal or discretized signal) and calculates the posterior distribution of states given the signal by Bayes rule, denoted as $\pi(\hat{\theta}|\hat{v})=\frac{\pi(\hat{\theta},\hat{v})}{\pi(\hat{v})}$. We pick the action with higher expected payoff under the posterior distribution on the current observation. We repeat this process for all observations and then take the expectation on all the rational benchmarks we get. We can take the conditional expectation across different conditions, e.g., different explanations. This calculation is illustrated in Algorithm 4 in Appendix A.
@@
(Approximating) Mis-reliant rational benchmark $\textsc{R}^{\text{m}}$. The mis-reliant rational benchmark is the expected score of a rational decision-maker with the same behavioral reliance level as the human participant. To calculate this, we simulate the rational decision-maker completing the same set of trials as the behavioral decision-makers do but additionally constrain the reliance level to be the same as the reliance level produced by the behavioral decision-makers. In our example experiment, each behavioral decision-maker completes $20$ trials with reliance levels, $\gamma^{b}=\Pr[a^{b}=y^{AI}|y^{AI}\neq y^{H}]$. As the rational decision-maker traverses the $4000$ observations, like behavioral participants she should engage in $20$ consecutive trials for each set. Suppose that the signals that the rational decision-maker receives in the 2020 consecutive trials are $\hat{v}_{1},\ldots,\hat{v}_{20}$. For each signal $\hat{v}_{i}$, the rational decision-maker knows the posterior payoffs, i.e., $\text{\bf E}_{\pi(\hat{\theta}|\hat{v}_{i})}\![{S(y^{AI},y)}]$ and $\text{\bf E}_{\pi(\hat{\theta}|\hat{v}_{i})}\![{S(y^{H},y)}]$. Then, the rational decision-maker ranks the signals in decreasing order of $\text{\bf E}_{\pi(\hat{\theta}|\hat{v}_{i})}\![{S(y^{AI},y)}]-\text{\bf E}_{\pi(\hat{\theta}|\hat{v}_{i})}\![{S(y^{H},y)}]$ and accepts the AI recommendation from the first signal in the sorted list, up to a $\gamma^{b}$ fraction of $20$ signals. We take the expectation over all observations (or conditionally on the manipulated variable of interest depending on the study design). This calculation is illustrated in Algorithm 6 in Appendix A. Note that estimation of the mis-reliant rational benchmark faces the same risk of overfitting as the rational benchmark. When the signal space is infinite, we approximate the mis-reliant rational benchmark the same way that we do the rational benchmark by calculating the upper- and lower-bound.